# An Analytic Characterization of the Limb Asymmetry - Transit Time Degeneracy

### v1

@@ 3.2 Connection to literature measurements
Our numerical method described in Section 3.1 resembles the method by which most, if not all, literature measurements of exoplanet transit times are determined. In reality, one obtains a transit observation of an exoplanet which may or may not have asymmetric limbs. The data are then typically fit using a uniform-limb model for the planet’s disk, using batman or similar light curve models, typically using a Bayesian parameter sampling technique, such as emcee. Therefore, using this numerical method we are able to directly compare our analytical predictions to what one may derive from a real observation.
@@ 4 Comparing our analytical and numerical results
First, we recorded the reference light curves generated for the “perfect” observation, determined the first (T1) and last (T4) contact times of each light curve, and calculated the change in both relative to the N = 0 (i.e. uniform-limb) case as N increased. Note the slight difference that this is comparing an increasing $R_{\mathrm{morning}}$ to the original $R_{\mathrm{evening}}$, and not comparing to a co-varying $R_{\mathrm{eff}}$, so only T1 will change. We compared these changes to what Equations 2.2 and 2 predict and find they perfectly agree, as shown in Figure 3. In terms of magnitude, we find that T1 occurs over 40 seconds earlier when the morning limb radius is increased by 5 scale heights. As a further test, we find the same behavior for T4 upon adjusting only the evening limb radius.
@@
Next, we computed the difference between the best-fit time of conjunction of the uniform-limb model for the “perfect” observation’s simulated data to the true value, $\Delta t_{c}$. This is shown as the circular points in Figure 4. We compared this numerical time offset to our analytical model’s prediction for $\Delta t_{c}$ (Equation 2), shown as the blue line. The numerical model finds a larger $\Delta t_{c}$ than the analytic model by a factor of about 10% at all $N$, reaching nearly a three second discrepancy at our maximal case of limb asymmetry of N = 5. This discrepancy consistently appeared, with approximately the same relative magnitude, in all of our numerical vs. analytical comparisons regardless of our specific choice of planetary parameters, limb darkening coefficients, and noise parameters. We discuss the underlying reason for this in the following section.
Figure 4: Measured times of conjunction compared to the true times of conjunction, as a function of the injected difference in radius between the planet’s evening and morning limbs. The synthetic data used in these fits use an asymmetric limb transit model. The circular points show fitting results from our numerical model, which fits using a uniform-limb transit model. The blue line shows our analytical model’s prediction for this bias in $t_{c}$ (Equation 2). The discrepancy between the numerical model and the analytical model is due to the “extra-area compensation effect”, described in Section 4.1. The green line shows the analytic model when including this compensation effect (Equation 12).
@@ 4.1 Extra-area compensation when fitting transits
We find that the discrepancy in time offsets results from the numeric model preferring larger values of $\Delta t_{c}$ which serve to minimize the residuals in the fits to our simulated data. This is a purely numerical effect, which we dub the “extra-area compensation effect”. We illustrate how the effect arises in Figure 5. In short, our analytic model solves for the $\Delta t_{c}$ that makes T1 and T4 match between the uniform-limb and asymmetric-limb planets, as shown in Panel C. However, due to the difference in size and shape of the disks, the asymmetric-limb planet still has “extra” occulting area, indicated by the blue arrows. This leads the asymmetric-limb planet to have a steeper ingress and egress compared to the uniform-limb planet, causing a large residual between their transit light curves despite having the same contact times. The numeric model compensates for this with an extra offset $\tau$ to the time of conjunction which has the effect of eliminating this “extra” area, shown in Panel D. In other words, the numerical model is drawn toward the flux-weighted mid-transit time.
@@
The planetary disks represented by our model light curves are shown for reference in panel A, with the uniform-limb planet’s disk in red and the asymmetric-limb planet’s disk in blue. Both disks have the same total area. Panels B, C, and D illustrate the planet approaching transit, with the star represented as the uniform yellow-orange disk. We start by treating the planets as having the same $t_{c}$, shown in panel B. Due to the limb asymmetry, fitting the uniform-limb model to an observation of the asymmetric-limb planet would lead to an offset $\Delta t_{c}$ in the inferred time of conjunction. This arises because the uniform disk’s time of first contact lags behind that of the asymmetric disk’s by time $\Delta t_{c}$. Our analytic model solves for this value of $\Delta t_{c}$ (Panel C). In our analytic model’s solution, however, the asymmetric disk has two regions of “extra” occulting area – due to its larger leading limb – which are indicated by the blue arrows in Panel C. Note that there is an equivalent “missing” area on the other side of the disk which affects egress. During the first half of ingress, this “extra” area leads to the asymmetric disk having a steeper ingress compared to the uniform disk, causing a large residual between the two light curves despite the two disks making first contact at the exact same time. Due to this residual, our numeric model disfavors this solution. The numeric model compensates for this “extra” area by offsetting the $t_{c}$ of the uniform disk by an extra amount $\tau$. As shown in Panel D, this has the effect of moving the edge of the uniform disk ahead of the asymmetric disk, so that the uniform disk now has its own “extra” area indicated by the red arrow. The preferred $\tau$ is that which makes the red and blue extra areas equivalent. The asymmetric disk’s steeper ingress is compensated by the uniform disk beginning ingress earlier, which has the net effect of minimizing the residual between their light curves. As a result, the “best-fit” solution from the numerical model is the situation shown in Panel D, where the uniform-limb model’s time of conjunction is offset by $\Delta t_{c}+\tau$.
@@
Figure 5: Schematic of the extra-area compensation problem. We generate a synthetic transit observation using the asymmetric-limb disk, shown in blue (Panel A). We fit these data using the uniform-limb disk shown in red. These two model planets are initially set to have the same time of conjunction $t_{c}$ (Panel B). Due to limb asymmetry imprinted into the data, this fit will return a biased $t_{c}$ when using this uniform-limb model. Panel C shows our analytic solution for this bias ($\Delta t_{c}$), which is the extra time the uniform-limb disk would travel until the two disks’ first and last contact times are equal. This value of $\Delta t_{c}$ is deemed a poor fit by the numeric model, however, as the “extra” occulting area of the asymmetric-limb disk in this arrangement leads to a large fitting residual. Rather, the numeric model prefers the solution shown in Panel D where the uniform-limb disk is translated even further ahead by time $\tau$. In this arrangement, there is now “extra” occulting area for the uniform-limb disk, indicated by the red arrow, which cancels out the asymmetric-limb disk’s extra area leading to a minimized residual, and thus a “better” fit.
@@
We can derive an analytical estimate for this compensation $\tau$ by comparing the path lengths travelled by the leading halves of each disk during ingress. Consider the alignment shown in Panel B of Figure 5, where the center of each disk is aligned. We position each disk halfway through ingress where the center of the disk makes contact with the edge of the stellar disk, as illustrated in Figure 6. We only show the leading half of each disk for simpler visualization. At this point, each half-disk must still travel some additional distance $d$ along $x$ before that half-disk is fully internally tangent (i.e., fully inside the stellar disk). For example, this additional distance for the asymmetric disk is given by the orange line at the top of Figure 6. The area of the planetary disk still remaining outside the stellar disk is analogous to the “extra” areas discussed above. In the coordinate system shown in Figure 6, the distance between the top of each planetary disk and the edge of the stellar disk is equal to {dmath} d = R_⋆- R_⋆^2 - y^2. Note that y also represents a length in this equation. When each disk has traversed approximately d/2, an equal area to what still remains outside the stellar disk will have entered within the stellar disk. This is analogous to the “extra” area compensation done by the numerical model, as shown in Panel D of Figure 5. Therefore, we can estimate $\tau$ from the difference in d/2 between the uniform-limb and asymmetric-limb disks, $\Delta d$/2. Based on Equation 5, this depends on the difference of their radii and the (fixed) radius of the star:
@@
Figure 6: Schematic showing the extra distance necessary for a disk hemisphere to travel between the point when the whole-disk center is contacting the disk and the point when the entire hemisphere will be internally tangent. The blue and red semi-circles represent the hemispheres of our uniform-limb and asymmetric-limb planets shown in Figures 1 and 6, and the orange circle is the star. This extra distance is equal to the x-value of the stellar disk at a given y, where the y-coordinate is described by the radius of the disk hemisphere. We use this as the basis for our derivation of the extra-area compensation given by Equation 12.
We computed the expected extra-area compensation for our “perfect” observation of WASP-39b scenario using Equation 13 and added it to our nominal analytical results shown in Figure 4. The result is shown as the green line in Figure 4 which agrees well with our numerical model. There is a small residual between the two, which slightly increases with increasing $N$. As alluded to previously, this is likely the result of us assuming the asymmetric-limb planet’s cross-section is composed of two connected semi-circles with discontinuity in the polar regions, as this discontinuity becomes more extreme with increasing $N$, among other assumptions discussed in Section 5.4. This residual is less than one-quarter of a second for the most extreme case of limb asymmetry. Achieving such precision on $t_{c}$ is beyond the capability of current instruments, and limb asymmetry of $N>5$ may be physically unrealistic.
@@
For each planet tested, the timing bias increases linearly from zero with increasing limb asymmetry. For modest (N = 1-2) amounts of limb asymmetry, the timing bias ranges from $\sim$1 - 10 seconds, which is small but detectable – especially with JWST observations. Note that in Figure 7, we give the absolute values for $\Delta t_{c}$. As mentioned in Section 2.2, the sign of $\Delta t_{c}$ depends on which limb is larger – $\Delta t_{c}>$ 0 when the evening limb is larger, and $\Delta t_{c}<$ 0 when the morning limb is larger. We find that the slope of each planet’s timing bias model differs and is correlated with the planet’s bulk scale height, meaning planets with larger scale height have a larger timing bias for a given value of $N$.
Figure 7: Differences between the measured times of conjunction compared to the true time of conjunction ($\Delta t_{c}$) as a function of the injected limb radius difference for simulated observations of representative exoplanets. Practically, these are the minimum a priori precision on $t_{c}$ necessary before limb asymmetry can be successfully measured from a transit observation. The solid lines show our analytical model’s predictions for $\Delta t_{c}$ using Equation 2. The points show our numerical model which mimic fitting real observations. We include the extra-area compensation effect in our analytic model using Equation 13. The key difference between the planets shown here is their atmospheric scale height. Higher scale height leads to larger $\Delta t_{c}$, meaning less a priori precision on $t_{c}$ is required to probe limb asymmetry.
@@ 5.1 Implications for investigating limb asymmetry
Error in a planet’s time of conjunction is a key systematic obstacle to investigating limb asymmetry via transmission spectroscopy. This raises the question of how accurately does a planet’s time of conjunction need to be known in order to reliably probe limb asymmetry on the planet. Answering this question was the primary motivation for our work, and our models for the timing bias (Section 2, and Figure 7) provide aid for future analyses in understanding and overcoming this obstacle.
@@
We note that our analysis has considered a transit observation using only a single filter, such as a photometric or band-integrated spectroscopic observation. Powell et al. (2019) showed the utility of instead using a multi-wavelength approach so that the spectral variation of limb asymmetry, when combined and forced to share a single $t_{c}$, actually works against itself to eliminate timing biases. Our results are complementary to such an approach. This multi-wavelength method will still lead to a best-fit $t_{c}$ with uncertainty, and our results demonstrate how well limb asymmetry may be explored with this remaining uncertainty.
Similarly, an alternate approach would be to simply measure the relative changes in $t_{c}$ as a function of wavelength. These chromatic variations, especially if aligned with spectral features of expected molecules in the planet’s atmosphere, would indicate the presence of limb asymmetry on the planet without needing an absolute $t_{c}$ for reference at all. While this approach indeed could easily identify the presence of a difference in atmospheric properties between limbs without concern of the degeneracy presented in this work, any deeper characterization of the limb-to-limb differences, for example fitting the transmission spectra of each limb across multiple instruments, would require an absolute $t_{c}$ value.
@@ 5.2 Implications of the extra-area compensation effect
In Section 4.1, we described an extra offset in $t_{c}$ measured for an asymmetric-limb planet due purely to numerical effects during the fitting process. The magnitude of this offset depends on the difference between the morning and evening radii (Equation 13, recalling the definition of $R_{\mathrm{eff}}$ from Equation 1). Therefore, per multiple of the scale height, this offset will also be larger for planets with larger scale heights as there is more physical occulting area to compensate for. With this in mind, we can also generalize Equation 13 by calculating $\tau$ as a function of $N$ for each of our example planets, which are linear functions, and fitting their slopes as a function of the scale height. We find that this extra-area compensation offset can therefore be approximated by
@@
Even in the most extreme cases, the magnitude of this offset is relatively small. For WASP-107b for example, which has one of the highest atmospheric scale heights of any exoplanet, the offset is on the order of one second or less for modest (N $<$ 2) limb asymmetry. Nevertheless, the existence of this extra-area compensation effect highlights the inherent challenge in accurately measuring a planet’s time of conjunction. Not only must potential astrophysical biases, such as the timing bias or dynamical transit timing variations (discussed in the next section), be considered, but biases introduced by numerical techniques in the analysis can have an impact as well. Fortunately, this particular numerical bias is relatively small – far smaller than the uncertainty of most current measurements achieved by Kepler, TESS, Spitzer, and ground-based observations – and will only exist if the planet indeed has limb asymmetry. Still, this illustrates the types of timing biases that must be newly considered when performing high-precision transit observations with JWST.

### v2

@@ 3.2 Connection to Literature Measurements
Our numerical method described in Section 3.1 resembles the method by which most, if not all, literature measurements of exoplanet transit times are determined. In reality, one obtains a transit observation of an exoplanet that may or may not have asymmetric limbs. The data are then typically fit using a uniform-limb model for the planet’s disk, using batman or similar light-curve models, typically using a Bayesian parameter sampling technique, such as emcee. Therefore, using this numerical method we are able to directly compare our analytical predictions to what one may derive from a real observation.
@@ 4 Comparing Our Analytical and Numerical Results
First, we recorded the reference light curves generated for the “perfect” observation, determined the first (T1) and last (T4) contact times of each light curve, and calculated the change in both relative to the $N$ = 0 (i.e. uniform-limb) case as $N$ increased. Note the slight difference that this is comparing an increasing $R_{\mathrm{morning}}$ to the original $R_{\mathrm{evening}}$, and not comparing to a covarying $R_{\mathrm{eff}}$, so only T1 will change. We compared these changes to what Equations 2.2 and 2 predict, and found they perfectly agree, as shown in Figure 3. In terms of magnitude, we found that T1 occurs over 40 s earlier when the morning limb radius is increased by five scale heights. As a further test, we found the same behavior for T4 upon adjusting only the evening limb radius.
@@
Next, we computed the difference between the best-fit time of conjunction of the uniform-limb model for the “perfect” observation’s simulated data to the true value, $\Delta t_{c}$. This is shown as the circular points in Figure 4. We compared this numerical time offset to our analytical model’s prediction for $\Delta t_{c}$ (Equation 2), shown as the blue line. The numerical model found a larger $\Delta t_{c}$ than the analytic model by a factor of about 10% at all $N$, reaching nearly a 3 second discrepancy at our maximal case of limb asymmetry of $N$ = 5. This discrepancy consistently appeared, with approximately the same relative magnitude, in all of our numerical vs. analytical comparisons regardless of our specific choice of planetary parameters, limb darkening coefficients, and noise parameters. We discuss the underlying reason for this in the following section.
Figure 4: Measured times of conjunction compared to the true times of conjunction, as a function of the injected difference in radius between the planet’s evening and morning limbs. The synthetic data used in these fits use an asymmetric-limb transit model. The circular points show fitting results from our numerical model, which fits using a uniform-limb transit model. The blue line shows our analytical model’s prediction for this bias in $t_{c}$ (Equation 2). The discrepancy between the numerical model and the analytical model is due to the “extra-area compensation effect,” described in Section 4.1. The green line shows the analytic model when including this compensation effect (Equation 12).
@@ 4.1 Extra-area Compensation When fitting transits
We find that the discrepancy in time offsets results from the numeric model preferring larger values of $\Delta t_{c}$ that serve to minimize the residuals in the fits to our simulated data. This is a purely numerical effect, which we dub the “extra-area compensation effect”. We illustrate how the effect arises in Figure 5. In short, our analytic model solves for the $\Delta t_{c}$ that makes T1 and T4 match between the uniform-limb and asymmetric-limb planets, as shown in Panel C. However, due to the difference in size and shape of the disks, the asymmetric-limb planet still has “extra” occulting area, indicated by the blue arrows. This leads the asymmetric-limb planet to have a steeper ingress and egress compared to the uniform-limb planet, causing a large residual between their transit light curves despite having the same contact times. The numeric model compensates for this with an extra offset $\tau$ to the time of conjunction, which has the effect of eliminating this “extra” area, shown in Panel D. In other words, the numerical model is drawn toward the flux-weighted mid-transit time.
@@
The planetary disks represented by our model light curves are shown for reference in panel A, with the uniform-limb planet’s disk in red and the asymmetric-limb planet’s disk in blue. Both disks have the same total area. Panels B, C, and D illustrate the planet approaching transit, with the star represented as the uniform yellow-orange disk. We start by treating the planets as having the same $t_{c}$, shown in panel B. Due to the limb asymmetry, fitting the uniform-limb model to an observation of the asymmetric-limb planet would lead to an offset $\Delta t_{c}$ in the inferred time of conjunction. This arises because the uniform disk’s time of first contact lags behind that of the asymmetric disk’s by time $\Delta t_{c}$. Our analytic model solves for this value of $\Delta t_{c}$ (Panel C). In our analytic model’s solution, however, the asymmetric disk has two regions of “extra” occulting area – due to its larger leading limb – which are indicated by the blue arrows in Panel C. Note that there is an equivalent “missing” area on the other side of the disk that affects egress. During the first half of ingress, this “extra” area leads to the asymmetric disk having a steeper ingress compared to the uniform disk, causing a large residual between the two light curves despite the two disks making first contact at the exact same time. Due to this residual, our numeric model disfavors this solution. The numeric model compensates for this “extra” area by offsetting the $t_{c}$ of the uniform disk by an extra amount $\tau$. As shown in Panel D, this has the effect of moving the edge of the uniform disk ahead of the asymmetric disk, so that the uniform disk now has its own “extra” area indicated by the red arrow. The preferred $\tau$ is that which makes the red and blue extra areas equivalent. The asymmetric disk’s steeper ingress is compensated by the uniform disk beginning ingress earlier, which has the net effect of minimizing the residual between their light curves. As a result, the “best-fit” solution from the numerical model is the situation shown in Panel D, where the uniform-limb model’s time of conjunction is offset by $\Delta t_{c}+\tau$.
@@
Figure 5: Schematic of the extra-area compensation problem. We generate a synthetic transit observation using the asymmetric-limb disk, shown in blue (Panel A). We fit these data using the uniform-limb disk shown in red. These two model planets are initially set to have the same time of conjunction $t_{c}$ (Panel B). Due to limb asymmetry imprinted into the data, this fit will return a biased $t_{c}$ when using this uniform-limb model. Panel (C) shows our analytic solution for this bias ($\Delta t_{c}$), which is the extra time the uniform-limb disk would travel until the two disks’ first and last contact times are equal. This value of $\Delta t_{c}$ is deemed a poor fit by the numeric model, however, as the “extra” occulting area of the asymmetric-limb disk in this arrangement leads to a large fitting residual. Rather, the numeric model prefers the solution shown in Panel (D) where the uniform-limb disk is translated even further ahead by time $\tau$. In this arrangement, there is now “extra” occulting area for the uniform-limb disk, indicated by the red arrow, which cancels out the asymmetric-limb disk’s extra area, leading to a minimized residual and thus a “better” fit.
@@
We can derive an analytical estimate for this compensation $\tau$ by comparing the path lengths traveled by the leading halves of each disk during ingress. Consider the alignment shown in Panel B of Figure 5, where the center of each disk is aligned. We position each disk halfway through ingress where the center of the disk makes contact with the edge of the stellar disk, as illustrated in Figure 6. We only show the leading half of each disk for simpler visualization. At this point, each half-disk must still travel some additional distance $d$ along $x$ before that half-disk is fully internally tangent (i.e., fully inside the stellar disk). For example, this additional distance for the asymmetric disk is given by the orange line at the top of Figure 6. The area of the planetary disk still remaining outside the stellar disk is analogous to the “extra” areas discussed above. In the coordinate system shown in Figure 6, the distance between the top of each planetary disk and the edge of the stellar disk is equal to {dmath} d = R_⋆- R_⋆^2 - y^2. Note that $y$ also represents a length in this equation. When each disk has traversed approximately $d$/2, an area equal to what still remains outside the stellar disk will have entered within the stellar disk. This is analogous to the “extra” area compensation done by the numerical model, as shown in Panel D of Figure 5. Therefore, we can estimate $\tau$ from the difference in $d$/2 between the uniform-limb and asymmetric-limb disks, $\Delta d$/2. Based on Equation 5, this depends on the difference of their radii and the (fixed) radius of the star:
@@
Figure 6: Schematic showing the extra distance necessary for a disk hemisphere to travel between the point when the whole-disk center is contacting the disk and the point when the entire hemisphere will be internally tangent. The blue and red semicircles represent the hemispheres of our uniform-limb and asymmetric-limb planets shown in Figures 1 and 6, and the orange circle is the star. This extra distance is equal to the $x$-value of the stellar disk at a given $y$, where the $y$-coordinate is described by the radius of the disk hemisphere. We use this as the basis for our derivation of the extra-area compensation given by Equation 12.
We computed the expected extra-area compensation for our “perfect” observation of WASP-39b scenario using Equation 13 and added it to our nominal analytical results shown in Figure 4. The result is shown as the green line in Figure 4 which agrees well with our numerical model. There is a small residual between the two, which slightly increases with increasing $N$. As alluded to previously, this is likely the result of us assuming the asymmetric-limb planet’s cross-section is composed of two connected semicircles with discontinuity in the polar regions, as this discontinuity becomes more extreme with increasing $N$, among other assumptions discussed in Section 5.4. This residual is less than one quarter of a second for the most extreme case of limb asymmetry. Achieving such precision on $t_{c}$ is beyond the capability of current instruments, and limb asymmetry of $N>5$ may be physically unrealistic.
@@
For each planet tested, the timing bias increases linearly from zero with increasing limb asymmetry. For modest ($N$ = 1-2) amounts of limb asymmetry, the timing bias ranges from $\sim$1 - 10 s, which is small but detectable – especially with JWST observations. Note that, in Figure 7, we give the absolute values for $\Delta t_{c}$. As mentioned in Section 2.2, the sign of $\Delta t_{c}$ depends on which limb is larger – $\Delta t_{c}>$ 0 when the evening limb is larger, and $\Delta t_{c}<$ 0 when the morning limb is larger. We find that the slope of each planet’s timing bias model differs and is correlated with the planet’s bulk scale height, meaning planets with larger scale height have a larger timing bias for a given value of $N$.
Figure 7: Differences between the measured times of conjunction compared to the true time of conjunction ($\Delta t_{c}$) as a function of the injected limb radius difference for simulated observations of representative exoplanets. In practice, these constitute the minimum a priori precision on $t_{c}$ necessary before limb asymmetry can be successfully measured from a transit observation. The solid lines show our analytical model’s predictions for $\Delta t_{c}$ using Equation 2. The points show our numerical model, which mimics fitting real observations. We include the extra-area compensation effect in our analytic model using Equation 13. The key difference between the planets shown here is their atmospheric scale height. Higher scale height leads to larger $\Delta t_{c}$, meaning less a priori precision on $t_{c}$ is required to probe limb asymmetry.
@@ 5.1 Implications for Investigating Limb Asymmetry
Error in a planet’s time of conjunction is a key systematic obstacle to investigating limb asymmetry via transmission spectroscopy. This raises the question of how accurately a planet’s time of conjunction needs to be known in order to reliably probe limb asymmetry on the planet. Answering this question was the primary motivation for our work, and our models for the timing bias (Section 2, and Figure 7) provide aid for future analyses in understanding and overcoming this obstacle.
@@
We note that our analysis has considered a transit observation using only a single filter, such as a photometric or band-integrated spectroscopic observation. Powell et al. (2019) showed the utility of instead using a multiwavelength approach so that the spectral variation of limb asymmetry, when combined and forced to share a single $t_{c}$, actually works against itself to eliminate timing biases. Our results are complementary to such an approach. This multiwavelength method will still lead to a best-fit $t_{c}$ with uncertainty, and our results demonstrate how well limb asymmetry may be explored with this remaining uncertainty.
Similarly, an alternate approach would be to simply measure the relative changes in $t_{c}$ as a function of wavelength. These chromatic variations, especially if aligned with spectral features of expected molecules in the planet’s atmosphere, would indicate the presence of limb asymmetry on the planet without needing an absolute $t_{c}$ for reference at all. While this approach indeed could easily identify the presence of a difference in atmospheric properties between limbs without concern of the degeneracy presented in this work, any deeper characterization of the limb-to-limb differences – for example, fitting the transmission spectra of each limb across multiple instruments – would require an absolute $t_{c}$ value.
@@ 5.2 Implications of the Extra-area Compensation Effect
In Section 4.1, we described an extra offset in $t_{c}$ measured for an asymmetric-limb planet due purely to numerical effects during the fitting process. The magnitude of this offset depends on the difference between the morning and evening radii (Equation 13, recalling the definition of $R_{\mathrm{eff}}$ from Equation 1). Therefore, per multiple of the scale height, this offset will also be larger for planets with larger scale heights, as there is more physical occulting area to compensate for. With this in mind, we can also generalize Equation 13 by calculating $\tau$ as a function of $N$ for each of our example planets, which are linear functions, and fitting their slopes as a function of the scale height. We find that this extra-area compensation offset can therefore be approximated by
@@
Even in the most extreme cases, the magnitude of this offset is relatively small. For WASP-107b, for example, which has one of the highest atmospheric scale heights of any exoplanet, the offset is on the order of one second or less for modest ($N$ $<$ 2) limb asymmetry. Nevertheless, the existence of this extra-area compensation effect highlights the inherent challenge in accurately measuring a planet’s time of conjunction. Not only must potential astrophysical biases, such as the timing bias or dynamical transit timing variations (discussed in the next section), be considered, but biases introduced by numerical techniques in the analysis can have an impact as well. Fortunately, this particular numerical bias is relatively small – far smaller than the uncertainty of most current measurements achieved by Kepler, TESS, Spitzer, and ground-based observations – and will only exist if the planet indeed has limb asymmetry. Still, this illustrates the types of timing biases that must be newly considered when performing high-precision transit observations with JWST.